OpenStudy (anonymous):

You saw a shirt for $97. You didn’t have the cash, so you borrowed $50 from your mum and $50 from your dad = $100. You bought the shirt, and had $3 change. You gave your dad $1 and your mum $1 and kept the other $1 for yourself. Now you owe your mum $49 and your dad $49. 49+49 = 98 + your $1 = 99. Where is the missing $1?

jimthompson5910 (jim_thompson5910):

There is no missing $1

jimthompson5910 (jim_thompson5910):

That $1 that you're pocketing goes to either parent, so if you gave say your mom the $1, then you would owe her 49 - 1 = 48 You would still owe your dad $49 So 49+48 = 97 Add in the change of $3 to get 97+3 = 100 So nothing is missing.